2 different areas contain duck nests. In area 1, there were 474 eggs of which 270 hatched. In area 2, there were 805 eggs of which 270 hatched. Find a 95% confidence interval for p1 – p2, the difference in the population means. Interpret the confidence interval in question 9. What does it say about the proportion of eggs that hatch in each of the two areas?
Answer #1

Here, , n1 = 474 , n2 = 805
p1cap = 0.5696 , p2cap = 0.3354


Standard Error, sigma(p1cap - p2cap),
SE = sqrt(p1cap * (1-p1cap)/n1 + p2cap * (1-p2cap)/n2)
SE = sqrt(0.5696 * (1-0.5696)/474 + 0.3354*(1-0.3354)/805)
SE = 0.0282

For 0.95 CI, z-value = 1.96
Confidence Interval,
CI = (p1cap - p2cap - z*SE, p1cap - p2cap + z*SE)
CI = (0.5696 - 0.3354 - 1.96*0.0282, 0.5696 - 0.3354 + 1.96*0.0282)
CI = (0.1789 , 0.2895)